Aluminium metal can be made from aluminium ions by electrolysis. What mass of aluminium metal would be recovered if a current of 2.50 × 103 A passed through the solution for 15.0 minutes? Assume the yield is 100%.
A. 210 g
B. 105 g
C. 70 g
D. 240 g
1
Expert's answer
2020-06-10T07:36:01-0400

We know that


"w = \frac{MIt}{nF}"

Where

M= Molar mass

I = current

t= time

n= charge on ion

F = Faraday's constant = 96500

So

"w=\frac{27\times2.5\times 10^3\times15\times 60}{3\times96500}\\=70gm"

Hence option (c)